OpenStudy (anonymous):

Explain how to do the following in good detail. Without a calculator, determine the exact value of each expression. Sin(pi/2) Cos(7pi/6) Tan(pi/6)

OpenStudy (dumbcow):

use a unit circle...radius of 1 cos refers to x distance from origin sin refers to y distance from origin pi/2 = 90 degrees, this is a quarter of a circle starting at 3 o'clock, go 90 degrees to left you are at 12 o'clock at this point you are vertical distance of 1 from origin --> sin(pi/2) = 1 7pi/6 = 7*180/6 = 210 degrees this is 180 + 30, so 30 degrees past half a circle at this point, both x,y are negative Now use geometry, a right triangle with hypotenuse of 1 and angles 30:60:90 side opposite 30 degree angle is 1/2 side opposite 60 degree angle is sqrt3/2 --> cos(7pi/6) = -sqrt3/2 pi/6 = 30 degrees we already did this above, except now x,y are positive sin(pi/6) = 1/2 cos(pi/6) = sqrt3/2 the tangent is sin/cos --> tan(pi/6) = (1/2)/(sqrt3/2) = 1/sqrt3 = sqrt3/3